OpenStudy (anonymous):

for what values of x is f(x) = 2x^3 -x^2 +2x concave up?

OpenStudy (solomonzelman):

you need to find \(\large\color{black}{f''(x) }\)

OpenStudy (solomonzelman):

can you tell me what the first derivative will be?

OpenStudy (solomonzelman):

(apply the power rule to first 2 terms, and the last term should not be a problem)

OpenStudy (solomonzelman):

@ilovereyvis_x3 , excuse me, are you there?

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

I was drinking water

OpenStudy (anonymous):

let me find the first derrivative

OpenStudy (solomonzelman):

go ahead.

OpenStudy (anonymous):

f'(x) = 6x^2 - 2x +2

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

now, the second derivative

OpenStudy (anonymous):

f''(x) = 12x - 2

OpenStudy (solomonzelman):

yes, right again

OpenStudy (solomonzelman):

When \(\large\color{black}{f''(x)>0 }\), the \(\large\color{black}{f(x) }\) is concave up.

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

Set: \(\large\color{blue}{0>12x - 2 }\) and give me the answer in interval notation that satisfies this blue statement

OpenStudy (anonymous):

1/6 > x ???

OpenStudy (solomonzelman):

yes, for any values of x that are greater than 1/6.

OpenStudy (solomonzelman):

So, it is concave up over the interval of \(\large\color{black}{\left(\begin{matrix} \frac{\LARGE 1}{\LARGE 6}, +\infty\\ \end{matrix}\right) }\)

OpenStudy (anonymous):

is that right ?

OpenStudy (solomonzelman):

you are done.

OpenStudy (solomonzelman):

yes, correct....

OpenStudy (anonymous):

thank you so much

OpenStudy (solomonzelman):

Not a problem!